how to solve sequence a(n)=12-3(n)
Answer by Edwin McCravy(20056)

how to solve sequence a(n)=12-3(n)
a(n)=12-3(n)

Substitute  n=1:  a(1) = 12-3(1)  = 12-3  = 9
Substitute  n=2:  a(2) = 12-3(2)  = 12-6  = 6
Substitute  n=3:  a(3) = 12-3(3)  = 12-9  = 3
Substitute  n=4:  a(4) = 12-3(4)  = 12-12 = 0
Substitute  n=5:  a(5) = 12-3(5)  = 12-15 = -3
Substitute  n=6:  a(6) = 12-3(6)  = 12-18 = -6
Substitute  n=7:  a(7) = 12-3(7)  = 12-21 = -9
Substitute  n=8:  a(8) = 12-3(8)  = 12-24 = -12
Substitute  n=9:  a(9) = 12-3(9)  = 12-27 = -15
Substitute n=10: a(10) = 12-3(10) = 12-30 = -18
Substitute ............ etc., etc., etc.,

So the sequence is 9, 6, 3, 0, -3, -6, -9, -12, 15, -18, etc., etc., etc,.

Edwin